Math problem - Gym members pay $3 per workout with a one time membership fee of $98. Nonmembers pay$10 per workout. How many workouts would both a member and a nonmember have to do to pay the same amount?
Answer by checkley77(12844)

98+3x=10x
3x-10x=-98
-7x=-98
x=-98/-7
x=14 workouts would equalize the costs.